The division-wide aptitude test in



Mathematics was conducted to students.



The mean of the test is 58 and the standard



deviation is 12. The scores also approximate



the normal distribution. What percent of the



scores is between 55 and 65?



A. 9.87%



B. 18.23%



C. 21.90%



D. 31.77%




1
Expert's answer
2022-03-30T08:50:25-0400

"P(55<X<65)=P(\frac{55-58}{12}<X<\frac{65-58}{12})=P(-0.25<X<0.58)=0.71904-0.40129=0.31775"